OpenStudy (anonymous):

The perimeter of a rectangle is 16 inches and its diagonal is sqrt(34) inches. Find its length and width. length= width=

OpenStudy (anonymous):

hello jahtoday!

OpenStudy (anonymous):

length = a & width = b

OpenStudy (anonymous):

hey!

OpenStudy (anonymous):

so 2a+2b=16 correct?

OpenStudy (anonymous):

and

OpenStudy (anonymous):

don't know the rest lol I am assuming pythagoreom

OpenStudy (anonymous):

yes very good the 2 side lengths are a and b and the hypotenuse is the diagonal of the rectangle

OpenStudy (anonymous):

cool so a^2+b^2=16

OpenStudy (anonymous):

then solve by system of solutions?

OpenStudy (anonymous):

wait whats the length of the diagonal?

OpenStudy (anonymous):

yeah mistype sqrt34

OpenStudy (anonymous):

ah because its sqrt the ^2 cancels it out

OpenStudy (anonymous):

but in Pythagorean the a^2 +b^2=c^2 in this case c is the diagonal. so yes

OpenStudy (anonymous):

thank u 4 the visual myiniaya

myininaya (myininaya):

:)

OpenStudy (anonymous):

so yes now that u have the right 2 equations u can solve by systems of solutions u mentioned before

OpenStudy (anonymous):

I am going the route of a=8-b on the right track?

OpenStudy (anonymous):

then substituting it in

myininaya (myininaya):

k looks good

OpenStudy (anonymous):

W=3, L=5

OpenStudy (anonymous):

Thank you Zak but working there

OpenStudy (anonymous):

I well come every one

OpenStudy (anonymous):

so substituting it into b^2+8-b=34

myininaya (myininaya):

L=3, W=5 or L=5, W=3

OpenStudy (anonymous):

changing into quadratic as b^2-b-26=0 but not thinking I am right

OpenStudy (anonymous):

b^2+(8-b)^2=34

OpenStudy (anonymous):

oh forgot the ^ thank you

myininaya (myininaya):

jah is right

OpenStudy (anonymous):

so should go to 2b^2-16b+30=0

myininaya (myininaya):

b^2+64-16b+b^2=34

OpenStudy (anonymous):

yup now u can factor or use the quadratic formula

OpenStudy (anonymous):

remember however that these are measurements and measurements have to b positive